OpenStudy (anonymous):

MEDAL AND FAN REWARD What value of x makes the denominator equal to zero? 2x+4 over 8x-12

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

8x-12 = 0 +12 +12 8x = 12 -- divide both sides by 8 x = 12/8 -- simplify x = 3/2

OpenStudy (anonymous):

a. -2 b. 0 c. 3/2 d. 12

OpenStudy (anonymous):

look at the work ^^

OpenStudy (anonymous):

oh sorry my bad i overlooked the answer part hahaha thanks so much for your work!

OpenStudy (anonymous):

No problem